OpenStudy (anonymous):

4.3 g of vitamin B-12 is dissolved in 74g of ether, (C2H5)2O at 25 degree Celsius. The change in vapor pressure amounts to 1.00 torr. Calculate the molar mass of this vitamin. Pressure of ether at 25 degree Celsius is 315 torr. Could someone please help me with this problem

OpenStudy (aaronq):

\(\Delta P_{solvent} =\chi_{solute}*P^o_{solvent}\) \(\chi_{solute}=\dfrac{n_{solute}}{n_T}=\dfrac{\dfrac{m_{solute}}{M_{solute}}}{\dfrac{m_{solute}}{M_{solute}}+n_{solvent}}\)